16 June 2009 UT - LBC Tech Night 2

Observers: O. Kuhn (mtn), J. Hill (Tucson, 1st half)
Operator: W. Wack
Instrument Support: J. Morris

Preparations

afternoon
  • John made some changes to the LBC code. See Issue 2214 for details. We found that last night, the mirror z in the headers was incorrect but the focus offset was correct. So the focus sequences and results were OK since they use FOCUSOFF. However, the absolute values of the focus offsets from last night are obsolete since we changed the Z22Zeros below.

  • I changed the rightchannelfocus.dat file to reflect the measurements from last night.

  • LBC gui does not work from lbtmu42 or from lbtmu06.

PM
  • incantation for dofpia to work from lbtmu06
    • source /home/lbcobs/LBCFPIA/lbcfpia/setup.csh
    • setenv LBTCONF_PATH /lbt/tcs/current.centos5/etc/lbt.conf

Open

Opened a bit late due to some minor telescope startup issues.

03:09 starting to focus/collimate

03:40 corrected pointing. Slew to SA107 (not photometric!!!) - observing this field simply to check focus on short exposures after making changes to every filter in the rightchannelfocus.dat file, and changing the LBC code.

04:10 Well... all of the images looked elongated in the same way - ellip ~0.1 and upper right to lower left. I let it run for a while but then it stopped with a red filter wheel #1 error - the filter did not reach position after 10 attempts. Running dofpia again.

04:31 stopping SA107 all red OB - I analyzed the first dither position - still some elongation upper right/lower left, but better. Seeing quite good now - 3pix fwhm - so going to check some filter focus offsets. No - instead going to check zeropoints for Z22.

Measuring Z22 Zeropoints

04:56 slew to field at ra=16, dec=25. Z22Zero = 2.05 for both blue and red. Dofpia... sequences: Red 050808--- 051408 and Blue 050755---051231. First image in Blue has point in middle (-ve sph) and first Red has donut (+ve sph)

05:21 changed BlueZ22Zero from 2.05 to 1.40 and RedZ22Zero from 2.05 to 3.05. Running dofpia again. Stop because see strong rings in Red pupils (which indicates the wrong sign of the correction - confirms the results of a few days ago).

Only just now are the primary mirrors settling into equilibrium.

05:34 Change BlueZ22Zero from 1.4 to 2.3 and RedZ22Zero from 3.05 from 1.95. Running dofpia again, followed by LongCoarse superfoc sequence: Red 054304..054705. Blue 054331...054809. Both still show the same trends as they did before - red from donut to point, blue from point to donut.

06:00 Change BlueZ22Zero from 2.3 to 2.5 - LC sequence is 060659...061137 and RedZ22Zero from 1.95 to 1.8 - LC sequence is 060629...061227.

06:20 Change BlueZ22Zero from 2.5 to 2.8 LC from 062917....063356 and RedZ22Zero from 1.8 to 1.7 LC from 062845...063354 Images from these sequences are attached to Issue #2213.

06:40 Change BlueZ22Zero from 2.8 to 2.95 and RedZ22Zero from 1.7 to 1.65 and dofpia again.

Based on the above sequences: we selected BlueZ22Zero = 2.95 and RedZ22Zero = 1.65 to use for the rest of the night. Admittedly the optimization is fairly soft. These values could certainly use some more tuning in the future, but they are vastly improved from last week.

Measuring Filter Focus Offsets

07:00 focus offsets:

filter focus fwhm file
dofpia (above)
r 0.056 2.74 focus1
V 0.025 2.70 fV1
i -0.0221 2.79 fi1
g +0.09 3.14 fg1
Y -0.046 2.60 fY1
U 0.315 3.33 fU1
dofpia
r 0.064 2.68 fr2
V 0.044 2.78 fV2
i -0.0055 2.52 fi2
g 0.098 3.11 fg2
Y -0.019 2.09 fY2
U 0.337 3.32 fU2
dofpia
i need to jiggle HP 2 controller
g 0.115 3.34 fg3

07:42 DX getting stuck - failing to reach position. Wayne slewed to zenith in preparation for going up to the cell to jiggle HP #2 controller. But at zenith, DX was focussing. So will slew to a star field at higher elevation - previous one was setting and had reached elev ~ 60 deg.

filter focus fwhm file
i -0.07 (limit !?)   fi4
g 0.107 3.18 fg4
Y -0.045 2.77 fY3
U 0.358 3.27 fU3
r 0.0 (limit!?) 3.08 fr3
V 0.0445 3.39 fV3
i -0.07 (limit again) 2.85 fi5
g 0.099 3.41 fg5

08:35 DX mirror not moving as expected - showing both X, Y limits (though X is red at about -1.5 not the -2.6 limit) and all other totals in yellow. Wayne forced it to panic - and initially it was not moving to the correct positions - platform positions not equal to total collimation. It seems to be doing a little better now as I am running dofpia. These are the new improved X-Y limits which are a 3 mm radius circle rather than a 2.6 mm halfwidth box.

filter focus fwhm file
i 0.007?? after eliminating almost all stars 2.54 fi6
g 0.095 good fit 2.72 fg6
r 0.05 2.55 again had to elimate all but one star - others had fwhm~60pxl! fr4
V 0.025 good fit 2.92 fV4

When I examine some r images with imexam, profile fits do not converge - give order 10's pix fwhm, so it is plausible that many stars must be eliminated.

filter focus fwhm file
dofpia
z -0.01 2.61 good fit fz1
B 0.061 3.05 fB1
I -0.0124 2.76 fI1
r(Blue) 0.100 3.37 frB1
R 0.0314 3.35 fR1
Us 0.424 4.30 fUs1

We had two images where it looked like DX had moved: 091539, first in i-SLOAN sequence and 095011, first of dofpia sequence.

10:10 A third trailed image caused problems in dofpia, so I stopped it., we cleared active optics. Finally we slewed to be at a higher elevation which has seemed to help DX a little. Now running dofpia again.

F972 0.13 (at lower limit - first point was severely comatic and deleted) 2.78 f1
no corresponding Blue sequence
z -0.0416 3.00 fz2
B consistent with 0.05 if last point disregarded - not a good fit 4.18 fB2
I -0.029 2.63 fI2
r(Blue) not a good fit   frB2
i 0.0113 2.94 fi6
g 0.122 3.76 fg6
dofpia

10:46 during this focussing, the first red pupil image again shows the streak - this is lbcr.20090616.104435.fits (other examples in issuetrak #2223

filter focus fwhm file
F972 0.179 0.83 f2
z have to still fit   fz3
B blue rotator not going to position 11:07 UT  

Close

11:22 04:22 - closing up. I stopped LBC systems and restarted and then stopped again. The Blue rotator seems to be moving as hoped - to the park position - as I shut down this second time. (No cmu reboot needed).

-- OlgaKuhn - 16 Jun 2009
I Attachment Action Size Date Who Comment
daq.20090616.gifgif daq.20090616.gif manage 51 K 08 Jul 2009 - 21:54 JohnHill Plot of Telescope and Mirror Temperatures
Topic revision: r4 - 08 Jul 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback